OpenStudy (anonymous):

If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system?

OpenStudy (anonymous):

the system is true

OpenStudy (tkhunny):

Is that true? Does it depend on ANY variable?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

thats all it says

OpenStudy (anonymous):

well if 4=4 then it is true

OpenStudy (tkhunny):

Hogwash. Please think more and type less. Is 4 = 4 a true statement?

OpenStudy (anonymous):

the system can't be solved

OpenStudy (tkhunny):

Does 4 = 4 depend on any value of x or y or z or t or s or q or p?

OpenStudy (anonymous):

the system is already solved @Catch.me

OpenStudy (camerondoherty):

lol

OpenStudy (anonymous):

Is it 4x = 4y or just 4 = 4?

OpenStudy (triciaal):

the values obtained satisfied both equations

OpenStudy (anonymous):

idk its on one of my assignments. so I should say its true?

OpenStudy (anonymous):

no 4 = 4 isn't a solution where x=5 or like that ? @Michiganboy21

OpenStudy (tkhunny):

You should answer my questions. This will lead you to the COMPLETE answer.

OpenStudy (anonymous):

\[ \begin{array}{lcl} x&=&x+1 \\ x\times 0&=&(x+1)\times 0\\ 0&=&0\\ 0+4&=&0+4\\ 4&=&4 \end{array} \]

geerky42 (geerky42):

I think what he mean is if we have like 2x+4y=8 and x+2y=4, then we would get 4=4 it cannot be solved. Because you can factor 2 out from 2x+4y=8 and you will have x+2y=4, basically, you are given two same equations with two unknown variables, so it cannot be solved.

geerky42 (geerky42):

So @Catch.me is correct

OpenStudy (tkhunny):

@catch.me Your argument is incorrect. There are infinitely many solutions. @JNoeller Your ONLY choices are 1) No solution or Cannot be Solved, 2) A single, unique solution, or 3) Infinitely many solutions.

OpenStudy (anonymous):

Getting to \(4=4\) is easy.

OpenStudy (tkhunny):

@wio Really? I think we have to assume that it was worked CORRECTLY in order to obtain 4 = 4.

OpenStudy (triciaal):

@JNoeller the question did not ask true or false. It ask what do you know about the system. It means there is a solution.

geerky42 (geerky42):

It cannot be solved because there is infinitely many possible value for variables, so there is no unique solution

OpenStudy (anonymous):

Well then what should I say on the paper?

OpenStudy (tkhunny):

The THREE possible outcomes are categorized as follows: 1) Something ridiculous, like 3 = 7, THEN there is no solution and it cannot be solved. 2) Something conditional, like x = 4 and y = 5, THEN there is a unique solution and it is so solved. 3) Tautology, like 4 = 4, THEN there are infinitely many solutions, since this result does not depend on anything conditional.

OpenStudy (anonymous):

yea you are right and for the same reason we don't divide by vector.@tkhunny

OpenStudy (tkhunny):

@geerky42 Your logic if flawed. Please read what you wrote. 1) It cannot be solved because 2) there is infinitely many possible value for variables, 3) so there is no unique solution Rephrase 1) No solution 2) Infinitely many solutions 3) No unique solution. Explanation: 1) Utterly disagrees with 2 2) This is correct. 3) Does not contradict #2

OpenStudy (triciaal):

Repost of @geerky42 Because you can factor 2 out from 2x+4y=8 and you will have x+2y=4, basically, you are given two same equations with two unknown variables, so it cannot be solved. this can be solved. it means that the values work in both equation and is an example of what i said, solution satisfies both equation

OpenStudy (tkhunny):

@JNoeller That was an awful lot of response in a very short time. Please read my posts carefully. These provide the complete solution and explanation. Hopefully, this will get everyone on the same page.

OpenStudy (tkhunny):

Given x + 4 = x + 4. What values for x satisfy this equation? Subtract x from both sides, giving 4 = 4. It is a tautology and there are infinitely many solutions.

geerky42 (geerky42):

There is difference between "no solution" and "cannot be solved" Say you have equation y=x and you are asked to solve it. do that make sense? no because there is infinitely many possible values for x and y that satisfy this equation. how can you "solve" y=x ?

OpenStudy (aum):

Example: 2x + 4y = 4 --- (1) and x + 2y = 2 ---- (2) They both represent the same line as the second one is gotten by dividing the first one by 2. From (2): x = 2 - 2y. Put this in (1): 2(2-2y) + 4y = 4 4 - 4y + 4y = 4 4 = 4. It means they are dependent equations. All points on the line are solutions. They have infinite number of solutions.

OpenStudy (anonymous):

okay thanks tkhunny

OpenStudy (anonymous):

Yes, it's infinite solutions, because it's coincidental lines. Hence the statement is true.

OpenStudy (tkhunny):

@geerky42 There is a bigger difference between "cannot be solved" and "infinitely many solutions". In the case of 4 = 4, "cannot be solved" is incorrect. It CAN be solved. There are infinitely many solutions. x = y is easily solved. Pick any pair (x,y), such that x = y and you will have a solution. If you mean a UNIQUE solution, you must state that. x = 2, y = 2 - Solved x = 3, y = 3 - Solved Unique solution? No.

geerky42 (geerky42):

alrights... you are right. excuse me for any inconveniences, I kind of have dyslexia what I trying to say is that there is no unique solution...

OpenStudy (tkhunny):

Very good. I hope we're on the same page. To explain a little more, give the result 4 = 4. If we pick x = 7, is 4 = 4 still true? Yes. Then x = 7 is a solution. If we pick x = 9, is 4 = 4 still true? Yes. Then x = 9 is a solution. If we pick x = -17, is 4 = 4 still true? Yes. Then x = -17 is a solution. The tautology 4 = 4 is independent of x (or any other variable expression). Pick anything you like and it will be a solution.

OpenStudy (tkhunny):

@preetha In 10 minutes, this generated 37 responses. I'd like to see another website do that!! :-)